OpenStudy (anonymous):

6r-7/10=r/4 show work please.

OpenStudy (anonymous):

cross multiply so that 4(6r-7)=10r Expand 24r-28=10r bring like terms together 24r-10r=28 14r=28 divide by 14... r=28/14 so... r=2

OpenStudy (anonymous):

im confused about the cross multiplying part, how do you do it and to what?

OpenStudy (anonymous):

oh nevermind , i get it. thanks .